OpenStudy (anonymous):

Two straight lines are perpendicular to each other. The sum of their slopes is equal to 1. What is the difference of their slopes ? Two straight lines are perpendicular to each other. The sum of their slopes is equal to 1. What is the difference of their slopes ? @Mathematics

OpenStudy (anonymous):

the difference of the slopes is \[\pm \sqrt{5}\]

OpenStudy (anonymous):

Thanks, Arnab.

OpenStudy (anonymous):

Is there a visual way of getting to the answer, than the quadratic ?

OpenStudy (anonymous):

yeah, of course. i am explaining

OpenStudy (anonymous):

take the slopes m1 and m2 then m1*m2=-1 m1+m2=1 so, m1-m2=\[\pm \sqrt{(m1+m2)^{2}-4m1m2}=\pm \sqrt{5}\]

OpenStudy (anonymous):

m1*m2=-1 M1+m2=1 . m1*m2=-1 (because perpendicular) m1+m2=1 (sum of gradients = 1) M1=-1/m2 M1+m2-1=0 m1+1/m2=0 simultaneous equations using subsitituion m1-m1+m2-1-1/m2=0 m2-1/m2-1=0 m2^2-m2/m2+1=0 (times through by m2 m2^2-1+1=0 M2=0 How did you get +-sqrt5?

OpenStudy (anonymous):

@payals, u got it?

OpenStudy (anonymous):

Yes, thanks. Would be even better if there is some visual intuition to it.

OpenStudy (anonymous):

@ullere, i cant get ur process

OpenStudy (anonymous):

maybe it will be better if u take help of equation button :)

OpenStudy (anonymous):

sorry it might be clearer if I use x and y x*y=-1 (perpendicular lines) x+y=1 (sum of lines = 1) first equation x*y=-1 (divde both sides by y X*y/y=-1/y, x=-1/y (add 1/y to both sides) x + 1/y = 0 second equation x+y=1 (subtract 1 from both sides) x+y-1=0 so now I use substitution to simultaneously solve the equations x+1/y = 0 x+y+1 = 0 x+y+1=x+1/y (subtract x+1/y from both sides x +y +1 - x - 1/y = 0 y+1 -1/y = 0 (multiply equation y) =y^2+y-1=0 b2-4ac = 5 yeah I must have messed up last time, your answer is right I reckon.